The starter motor of a car engine draws a current of 200 A from the battery. The copper wire to the motor is 5.40 mm in diameter and 1.2 m long. The starter motor runs for 0.660 s until the car engine starts.

How far does an electron travel along the wire while the starter motor is on?
Answer #1

(a) current is amount of charge/time => amt of charge = current x time=200A x 0.66 sec

= 132 C

(b) We need to find the drift velocity of electrons for this part. The current I = nev_dA, => v_d = I/neA

or v_d = 200A/(pi(0.0054/2)^2 *8.5*10^28*1.6*10^-19 = 6.42*10^-4 m/s

. The distance travelled by electrons in 0.66 sec

= v_d x (0.66 sec)

= 4.24*10^-4 m